OpenStudy (anonymous):

Can someone please help me with this question on my Algebra quiz? I have literally done this problem a million times and not getting any of the right answers. (its not multiple choice) HELP ASAP . thank you Find 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2.

OpenStudy (anonymous):

By the way the the numbers after the variables are the exponents, so "2x2= 2 to the second power."

OpenStudy (anonymous):

so 2x^2 -2z^4 +2y -x^2 +z^4 if: x = -4, y = 3, and z = 2. Plug in the values 2(-4)^2 - 2(2)^4 +2(3) -(-4)^4 +(2)^4 take care of exponent then multiple by coefficient. I'll do first one: 2(-4)^2 = (-4*-4)*2 = 32

OpenStudy (anonymous):

What is 2(2)^4?